OpenStudy (zzelinski):

4/5 y - 1/2 y = 3 help please!!!

OpenStudy (zzelinski):

\[\frac{ 4 }{ 5 } y - \frac{ 1 }{ 2 } y = 3\]

OpenStudy (zzelinski):

\[-\frac{ 7 }{ 2 } \div \frac{ 3 }{ 5 }\] I don't remember how to do this....

OpenStudy (jim766):

multiply everything by a common denominator to get rid of the fractions. In this case 10 \[10\frac{ 4 }{ 5 }y - (10)\frac{ 1 }{ 2 }y= 10(3)\] \[\frac{ 40 }{ 5 }y-\frac{ 10 }{ 2 }y=30\] now simplify and solve

OpenStudy (jim766):

on the 2nd one change it to multiplication by taking the reciprocal of the 2nd fraction \[\frac{ -7 }{ 2 }\div \frac{ 3 }{ 5 }=\frac{ -7 }{ 2 }\times \frac{ 5 }{ 3 }\] to multiply fractions....just multiply straight across -7 * 5 ------ = ------ 2*3

OpenStudy (jim766):

make sense?

OpenStudy (zzelinski):

The first one doesn't, but the second one does

OpenStudy (jim766):

40/5y - 10/2 y = 30 8y - 5y = 30 3y = 30 y = 10

OpenStudy (zzelinski):

ohhh now I get it thanks!

OpenStudy (jim766):

yw